You get 5.3x10^12 electrons on your pants as you slid across a car seat. What would this charge be in Coulombs?
Answer #1

1 electron = 1.6*10-19 Coulombs

5.3*1012 electrons = 5.3*1012*1.6*10-19 C

= 8.48*10-7 Coulombs